For each of the models listed below, predict y when x 3. a) y 1.6 0.4log x b) log y 1.6+0.4x c) y 1.6+0.4vx d) y 1.6 (0.4* e) y 0.4x1.6x+ 3 a) y 1.79 (Round to two decimal places as needed.) b)y 2.8 (Round to two decimal places as needed.)
Answer #1

a) Log 3 = 0.47712

Therefore y = 1.6 + [0.4*0.47712) = 1.6 + 0.19 = 1.79

b) log y = 1.6 + 0.4*3 = 1.6 + 1.2 = 2.8. Therefore y = log-12.8 = 630.96

c) y = 1.6 + 0.4*sqrt(3) = 1.6 + 0.4*1.732 = 1.6 + 0.6928 = 2.2928 \approx   2.29

d) y = 1.6*(0.43) = 1.6* 0.064 = 0.1024 \approx0.10

e) y = 0.4 * 32 + 1.6*3 + 3 = 3.6 + 4.8 + 3 = 11.40